My job right now is to characterize the green PDH loops on each arm. Today, Jenne took me around and pointed at the optics and electronics involved. She then showed me how to lock the green beams to the arms (i.e. opening the shutters until you hit a TM00 shape on the transmitted beam camera). Before lunch, the y arm was easiest to lock, and the transmitted power registered at around 0.75.
After lunch, I took a laptop and SR785 down to the y end station. I unhooked the PDH electronics and took a TF of the servo (without its boost engaged, which is how it is currently running) and noise spectrum with the servo input terminated.
I then set up things a la ELOG 8817 to try and measure the OLTF. However, at this point, getting the beam to lock on a TM00 (or something that looked like it) was kind of tough. Also, the transmitted power was quite a bit less than earlier (~0.35ish), and some higher order modes were higher than that (~0.5). Then, when I would turn on the SR785 excitation, lock would be lost shortly into the measurement, and the data that was collected looked like nonsense. Later, Koji noted that intermittent model timeouts were moving the suspensions, thus breaking the lock.
We then tried to lock the x arm green, to little success. Koji came to the conclusion that the green input pointing was not very good, as the TM00 would flash much less brightly than some of the much higher order modes.
Tomorrow, I will measure the x arm OLTF, as it doesn't face the same timeout issue that is affecting the y arm.